Ex2: A very myopic person has a near point of 8 cm, and a far point of 20 cm. What power lens does he need to see clearly at a distance? What is his new near point with this vision correction?
Answer #1

s = ∞
s' = -20 cm

1/f = 1/s + 1/s' = 0 + 1/(-20)

f = -20 cm = -0.20 m

P = 1.(-.20) = -5 diopters

His new near point with this vision correction = infinity